Scoresby Inc. tracks the number of units purchased and sold throughout each year but applies its inventory costing method at the end of the year, as if it uses a periodic inventory system. Assume its accounting records provided the following information at the end of the annual accounting period, December 31.

Transactions Units Unit Cost
  a. Inventory, Beginning 1,500 $ 30
  For the year:
  b. Purchase, March 5 7,500 31
  c. Purchase, September 19 3,500 33
  d. Sale, April 15 (sold for $75 per unit) 2,400
  e. Sale, October 31 (sold for $78 per unit) 6,500
  f. Operating expenses (excluding income tax expense), $406,000
Required:
1. Calculate the number and cost of goods available for sale.
2. Calculate the number of units in ending inventory.
3.

Compute the cost of ending inventory and cost of goods sold under (a) FIFO, (b) LIFO, and (c) weighted average cost.

4.

Prepare an income statement that shows the FIFO method, LIFO method and weighted average method.
Answer #1

1) Calculate the number and cost of goods available for sale

Units Unit Cost Total Cost
Beginning inventory 1500 30 45000
Purchase 1 7500 31 232500
Purchase 2 3500 33 115500
Total 12500 393000

Number of goods available for sale = 12500

Cost of goods available for sale = 393000

2) Ending inventory = 12500-8900 = 3600 Units

3) Calculate ending inventory and cost of goods sold

FIFO LIFO Weighted average
ending inventory (3500*33+100*31) = 118600 (1500*30+2100*31) = 110100 393000/12500*3600 = 113184
Cost of goods sold 393000-118600 = 274400 393000-110100 = 282900 393000-113184 = 279816

4) Income statement

FIFO LIFO Weighted average
Sales 687000 687000 687000
Cost of goods sold 274400 282900 279816
Gross profit 412600 40410 407184
Operating expense 406000 406000 406000
Income from operation 6600 -1900 1184